# What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality

### v1

@@ 2.3 Sensor-to-Subjective Mapping Framework
To structure our investigation, we introduce the Sensor-to-Subjective Mapping Framework (S2S) in Figure 1, which conceptualizes how collaborative group behavior in MR can be interpreted across three perspectives: observable system-level interaction traces (sensor-based indicators), emergent outcomes from interaction (task performance metrics), and post-task reflective accounts from users (subjective perceptions). This framework does not assume that one perspective directly determines the other; it treats each as a complementary lens that captures different facets of collaboration.
@@ 2.3.1 From Signals to Perception
The first relationship examines whether behavioral signals captured through sensors reflect users’ internal interpretations of collaboration. We refer to this relationship as reflective interaction mapping. For instance, we ask whether shared gaze events, as measured by eye-tracking, correspond to a user’s sense that the group had shared attention. Similarly, do balanced patterns in voice activity align with reported conversational support in the post-hoc survey? This connection motivates our first set of hypotheses, which examine whether subjective perceptions of collaboration are reflected in the observable sensor-based measures.
@@ 2.3.3 Subjective Perception to Task Performance
Finally, we investigate whether users’ internal experiences correspond to observable group-level performance metrics. We refer to this relationship as a subjective influence on performance. While individual states, such as presence and cognitive load, are not directly measured through sensors, they shape how users engage with others and respond to collaborative demands. For example, users who report lower cognitive load may have been more available to coordinate or contribute [13], which could result in more balanced participation and faster task completion. This mapping supports hypotheses that investigate whether the way users feel about their own experience or the group as a whole is meaningfully associated with collaborative outcomes.
@@ 2.4 Hypotheses
H3.1: Higher individual perceived presence and perception of contribution is reflected in more frequent shared attention events among group members.
@@ 3.2 Materials
To capture sensor data on user interactions, we used the Meta Quest Pro [32]. The integrated sensors enabled audio recording, eye-tracking, six-degrees-of-freedom (6 DoF) simultaneous localization and mapping (SLAM) tracking, and MR capabilities. The collaborative application is developed using Unity with Meta XR packages [31]. Unity and Meta XR API’s are used in custom scripts to record the audio, location, gaze-object intersections, and user interactions with virtual objects. For precise manipulation of virtual objects, we use the controller integrated with the Meta Quest Pro [32]. The interaction recording and tracking is validated by others [43]. Participants were invited to a shared lab room with a designated space cleared of materials to minimize distractions. They were informed that they could move freely within this area during the task.
@@
| Gender | 36 Male, 8 Female |
| Frequency of AR Experience | 21 Never Used, 14 Rarely, 6 Occasionally, 1 Frequently, 2 Frequently in the past |
| Frequency of VR Experience | 19 Never Used, 13 Rarely, 7 Occasionally, 4 Frequently, 1 Frequently in the past |
| Frequency of Gaming | 4 Never Used, 8 Rarely, 14 Occasionally, 15 Frequently, 3 Frequently in the past |
| Familiarity to Other Members | 25 No Members, 11 One Member, 6 Two Members, 2 Three Members |
@@ 3.3 Experimental Task
This section summarizes the cue, interaction, and feedback of our collaborative image-sorting task. The study was deliberately designed around a single collaborative task under one shared condition, without varying levels of network stressors or task types. Our goal was not to compare multiple experimental conditions but to examine the richness of group behavior in a naturalistic, unconstrained collaborative setting. Prior work has shown that tightly controlled comparisons can obscure the variability and fluidity of real-time collaboration in MR environments [65]. Instead, we focused on collecting high-resolution behavioral data and subjective reflections during a consistent, shared experience across groups. Each group of four participants completed one image-sorting task using the same images and categories. Participants were instructed to work together to reach a consensus on the grouping of each image.
@@ 3.3.1 Primary Task
Participants were tasked with sorting 28 carefully selected images from the Open Affective (OASIS) dataset [25]. The OASIS dataset comprises 900 validated images by over 822 participants for pleasantness and arousal ratings. The images were selected to represent a range of emotions while excluding potentially distressing content such as violence or graphic imagery. Participants sorted the selected images into one of six emotion categories randomly chosen from Russell’s circumplex model of affect [45]. These categories included angry, bored, relaxed, tense, pleased, and frustrated. Image sorting tasks have been shown to foster decision-making, communication, and social coordination by building shared mental models and group alignment [8]. We focus on this collaborative task marked by asynchronous, flexible participation, where participants can contribute and modify inputs independently. This repeated image-sorting task, involving open-ended discussions on the emotions elicited by each image, allows us to observe a dynamic, iterative collaborative process among the group, where participants achieve a shared goal through incremental steps and consensus. Each group of four participants completed one image-sorting task on the same 28 images and categories. Participants were instructed to work together as a group to reach a consensus on the label of each image, with no time limit for completing the task, allowing participants to engage in deliberate discussion and negotiation. The labels are not placed in a fixed position, allowing participants to organize and use the room space however they want. The task ends once they inform the researcher that they agree with the image groupings.
@@ 3.3.2 Virtual Scene and Cues
As shown in What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality (left, headset’s first-person perspective), all 28 virtual images are scattered around the room, and all six emotion categories labels (gray virtual plates) are pasted a little higher than where the images are scattered. Participants can view these images and labels at all times via their headsets. For participants, the cue to start the interaction is not defined by the researcher but decided by each participant which image they want to discuss with the group to sort. This lack of structure in cues is by design, as our goal is to observe open and free collaborations without participants taking turns or being directed by the flow of the virtual scene designed by the researcher. The participants are assumed to take the cue for virtual interaction from the other three group members as shown in What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality (right), where all four participants are physically close, probably examining the same image and discussing the final label.
@@ 3.3.3 Interaction and Feedback
To sort an image, the participants were asked to physically move the virtual image near the virtual label. Once the image is pasted close to the label, the image is recorded as sorted. Participants used a point-and-drag near-interaction motion with the grip buttons on their left or right controllers to move an image to the corresponding label. They pointed the controller at an image, pressed and held the grip button, and moved the image by guiding the controller. The image followed the controller’s pointer until the grip button was released, locking the image in its final position. Participants could only grab objects within reach and were instructed to release the grip button to secure an image in place, as demonstrated in Figure 2. This interaction closely mimics the physical action of picking up and placing objects. Once the image was positioned next to the label, it remained stationary when the grip button was released, providing feedback that the image was placed as intended. Only one participant could move each object at a time, but they were all allowed to hold different image objects simultaneously.
@@ 3.4 Measurements
This section outlines the measures we used to capture group behavior in the image sorting task, categorized into sensor level, task-related performance metrics, and subjective measures. At the sensor level we collected data from the headset using custom scripts. We recorded the audio signal from the microphone, positions for location, and eye-tracking data to capture the data for conversation, proximity, and shared attention as described in §2.1.2.
@@
Finally, we collect several subjective measures through post-exposure surveys after the image sorting task, such as presence (subjective feeling of being present in a virtual environment) with the IPQ [47] and PQ [62] questionnaires, cognitive load through NASA-Task Load Index (NASA-TLX) [20]222We are aware of the criticism surrounding the use of NASA-TLX. We applied it to measure “perceived” cognitive workload, rather than actual mental load [30]., and perception of the group behavior through a custom-designed questionnaire. The PQ evaluates factors such as the possibility to act and examine, realism, self-evaluation, and interface quality. The IPQ measures spatial and general presence, realism, and involvement. The presence scores are derived from 33 items (14 IPQ and 19 PQ, the cognitive load score is derived from 5 items from NASA-TLX, and group behavior from 5 items from our custom-designed survey on a 7-point scale. We developed a custom questionnaire to measure participants’ perspectives of how their group interacted, as shown in Table 2. The proposed group behavior characterization questionnaire assesses key aspects such as contribution awareness, shared attention, proximity impact, conversational support, and overall group collaboration. Each question is informed by established research to ensure relevance to our study’s context.
@@ 3.5 Study Procedure
The procedure for the study involved several steps, as shown in Figure 3. Upon arrival at the laboratory, participants were given a detailed study information sheet about the study procedures, data collection, and privacy measures. The administering researcher also verbally briefed them on this information. They were also verbally instructed on the headset interactions and controller gestures needed to complete those interactions required during the tasks and how to perform them. The briefing included details about the visual stimuli used in the experiment, such as their color, shape, duration, cues, and feedback mechanism for the image-sorting task. Participants were given ample time to consider their participation in the study and were asked for their verbal consent. Following the briefing, participants were asked to fill out a demographic questionnaire, with questions including their gender, age, familiarity with technology, and with other group members.
MR headsets were then distributed to the participants, and they were instructed to calibrate the focus and fit of the headset for maximum comfort. Before the main task, participants completed a tutorial application with two images and two categories not included in the main task to prevent learning effects. This tutorial task aimed to familiarize them with the task interactions in terms of gestures and get comfortable with using the point-and-drag interaction from the controller. Following this, the group proceeded with the main image sorting task, where they were tasked to sort images into six different categories. They were informed that there was no time limit for the task and that the main requirement was for them to reach a consensus on the image sorting for the task to end. To encourage a more natural and unconstrained group behavior, participants were not informed that they were being timed or evaluated on their accuracy.
Upon task completion, a post-task questionnaire assessing their cognitive load, presence, and perspective of what they think about how the collaboration went with the other group members in the task condition. The time it took for each group to complete the task varied; however, in general, the total duration of the session, including consent, briefing, training, headset calibration, experiment, and questionnaires, took less than an hour.
4 Results
4.1 Perceived Group Behavior & Tasks Summary
We summarize the descriptive statistics illustrating participants’ subjective experiences with response variability in Table 3. Metrics such as PQ () and IPQ () indicate moderate presence levels. The PQ-REAL subscale has a mean of 5.16 (), suggesting moderate realism, while the ACT subscale shows a high mean of 5.8, indicating strong perceived action capability. INV subscale variability () highlights differing engagement levels. SP () reflects strong spatial awareness. The mean NASA TLX score of 2.3 points to a low perceived workload. Group cohesion and collaboration scored high ( and ), while group proximity varied (, higher SD), indicating differing closeness perceptions.
Metrics: Mean (), Standard Deviation (), Standard Error (SE), 95% Confidence Interval (CI), 5th/95th Percentiles (P5/P95), Min, and Max. The sub-scales are realism (PQ-REAL, IPQ-REAL), possibility to act (ACT), interface quality (IFQUAL), possibility to examine (EXAM), self-evaluation of performance (EVAL), involvement (INV), general presence (GP), and spatial presence (SP).
@@ 4.1 Perceived Group Behavior & Tasks Summary
Next, we present two complementary sets of group-level metrics. A summary of behavioral interaction statistics captured through system instrumentation is seen at Table 4. These include the number of images grabbed per participant, total grabs, label overrides, label changes, and task completion time. These metrics reflect the group’s engagement with the task, coordination complexity, and task performance. For example, Group 8 had a notably high number of total grabs (436) and the longest task duration (994.05s), suggesting prolonged deliberation or difficulty reaching consensus. In contrast, Group 5 completed the task most quickly (209.26s), with lower override activity, possibly reflecting more streamlined decision-making or higher initial agreement.
@@ 4.2 Collaboration via Sensor-based Indicators
This section discusses the participants’ reflections on collaboration, alongside behavioral indicators obtained from sensors and system logs, at both individual and group scales. Findings for shared gaze overlap frequency (indicating visual focus on the same virtual image), proximity duration (measuring time spent in physical proximity during interaction), and speaking proportion and variance (representing turn-taking and conversational balance via total speaking time, frequency, and variance in participation) can be seen in Table 6. Normality checks for subjective collaboration scores were conducted using the Shapiro-Wilk test ( at the group level; - at the individual level), determining choice of Spearman or Pearson test based on sensor metric normality. Highlighted table rows indicate statistically significant correlations (), offering quantitative insight into how subjective assessments correlate with interaction behavior patterns.
@@ 4.4 Experience via Sensor-based Indicators
We first present the correlations between presence and cognitive load (NASA-TLX) scores and sensor-derived indicators of shared attention, proximity, and interaction balance in Table 8. Sensor-based metrics were drawn from the same behavioral indicators introduced earlier, capturing visual coordination, physical closeness, and contribution variability. Presence was not significantly associated with shared attention metrics (shared attention percentage: , ), proximity (mean pairwise distance: , ), or interaction balance (variance in image grabs: , ). This suggests that the individual sense of presence was not directly mirrored in the observable group interaction patterns. For cognitive load, however, we found a positive association with shared gaze overlap at the individual level (, ), indicating that users reporting higher mental effort tended to participate in more episodes of shared attention. Other indicators, including proximity and interaction balance, did not show significant relationships with cognitive load (, ). These results suggest that internal experience is partially reflected in the observable sensor-based metrics of group behavior. In particular, shared visual focus may demand or reinforce cognitive effort. At the same time, spatial positioning and contribution balance may reflect other aspects of collaboration beyond individual mental workload or presence.
@@ 5 Discussion
This section analyzes findings in the context of our S2S framework by assessing behavioral indicators from sensors and system logs, subjective experiences, and group reflections. We evaluate how these signals match participants’ reported collaboration experiences and task results, revealing insights into how collaboration is both enacted and experienced. Significant associations were identified between sensed group behavior and reported experiences. We distinguish between individual and group insights to evaluate the alignment’s strength and origin.
@@
Shared gaze frequency was a consistent indicator of perceived collaboration and shared attention at both analysis levels. Group-level correlations were particularly robust, highlighting shared visual focus as both an individual and collective phenomenon. In contrast, conversational support exhibited weaker correlations. A significant relationship between speaking balance and perceived support was only identified at the group level, implying that individual perceptions may not align with supportive speaking dynamics, but group-level participation does indicate communicative balance. We affirm H1 and H1.1, as multiple gaze-based features correlated with perceived collaboration and shared attention. We partially confirm H1.2, as conversational dynamics were significant solely in group-level aggregates.
@@
Behavioral indicators, particularly shared gaze frequency and proximity duration, demonstrated strong predictive ability for task completion time, especially at the group level. Groups that engaged in mutual observation and remained close completed tasks more rapidly, underscoring the importance of co-orientation in effective collaboration, as shown by our study’s metrics. Interaction balance, such as variance in image grabs or label overrides, was also linked with efficiency. Notably, these patterns persisted at both individual and group scales, reinforcing the role of participation equity in enhancing fluency. We accept H2, H2.1, and H2.2. Group-level behavior metrics, notably shared attention and proximity, predicted performance, and interaction balance indicated collaborative fluency across both levels.
@@
Sensory behavioral features were weakly associated with subjective states such as presence or cognitive load. A near-significant individual-level link was observed between cognitive load and shared gaze frequency, but it lacked strength and consistency at the group level. Presence did not correspond clearly with proximity or gaze patterns, suggesting a lack of visible manifestation of internal states in this context. Likewise, no significant relationship emerged between cognitive load and participation balance. H3 is partially accepted due to some individual-level signal, but not consistently. We reject H3.1, H3.2, and H3.3 due to a lack of strong evidence connecting presence or cognitive load to group-level behaviors.

### v2

@@ 2.3 Sensor-to-Subjective Mapping Framework
We introduce the Sensor-to-Subjective Mapping Framework (S2S) (Figure 1) to structure our investigation. It conceptualizes collaborative group behavior in MR across three perspectives: observable system-level interaction traces (sensor-based indicators), emergent outcomes from interaction (task performance metrics), and user-reported post-task reflections (subjective perceptions). Unlike prior work [4, 9, 48, 60], assuming one view determines another, our framework treats these perspectives as complementary, each capturing different facets of collaboration as described in §1.2.
@@ 2.3.1 From Signals to Perception
We first examine whether sensor-derived behavioral signals correspond to how users interpret collaboration. We call this a relationship. We call this relationship reflective interaction mapping. While subjective reports are post-hoc and summarized, we analyze objective indicators over task-level windows that match the granularity of users’ reflections. We do not compare specific events to single survey items but assess whether aggregate behavioral patterns (gaze overlap, turn-taking) are reflected in users’ overall experience. This avoids false precision and treats subjective reports as interpretive anchors rather than timestamped outcomes. For example, we ask whether shared gaze events align with a user’s sense of shared attention. Do balanced speech patterns correspond to reported conversational support? This connection motivates our first set of hypotheses, which examine whether subjective perceptions of collaboration are reflected in the observable sensor-based measures.
@@ 2.3.3 Subjective Perception to Task Performance
Finally, we investigate whether users’ internal experiences correspond to observable group-level performance metrics. We refer to this relationship as a subjective influence on performance. Though not sensor-measured, states like presence and cognitive load shape how users engage with others and respond to collaborative demands. For example, lower cognitive load may support better coordination [15], leading to more balanced participation and faster task completion. This mapping tests hypotheses that investigate whether subjective experiences (own experience or the group as a whole) align with observable group outcomes.
@@ 2.4 Hypotheses
H3.1: Higher individual perceived presence and perception of contribution are reflected in more frequent shared attention events among group members.
@@ 3.2 Materials
To capture sensor data on user interactions, we used the Meta Quest Pro [37]. The integrated sensors enabled audio recording, eye-tracking, six-degrees-of-freedom (6 DoF) simultaneous localization and mapping (SLAM) tracking, and MR capabilities. The collaborative app is built in Unity using Meta XR APIs to log audio, gaze, location, and object interactions [36]. For precise manipulation of virtual objects, we use the controller integrated with the Meta Quest Pro [37]. The interaction recording and tracking are validated by others [46]. Participants were invited to a shared lab room with a designated space cleared of materials to minimize distractions. They were informed that they could move freely within this area during the task.
@@
| Gender | 36 Male, 12 Female |
| Frequency of AR Experience | 22 Never Used, 15 Rarely, 6 Occasionally, 3 Frequently, 2 Frequently in the past |
| Frequency of VR Experience | 21 Never Used, 13 Rarely, 7 Occasionally, 6 Frequently, 1 Frequently in the past |
| Frequency of Gaming | 4 Never Used, 10 Rarely, 15 Occasionally, 16 Frequently, 3 Frequently in the past |
| Familiarity to Other Members | 25 No Members, 15 One Member, 6 Two Members, 2 Three Members |
@@ 3.3 Experimental Task
This section summarizes the cues, interaction, and feedback of our collaborative image-sorting task. The study was deliberately designed around a single collaborative task under one shared condition, without varying levels of network stressors or task types. Our goal was not to compare multiple experimental conditions but to examine the richness of group behavior in a naturalistic, unconstrained collaborative setting. Prior work has shown that tightly controlled comparisons can obscure the variability and fluidity of real-time collaboration in MR environments [70]. Instead, we focused on collecting high-resolution behavioral data and subjective reflections during a consistent, shared experience across groups. Each group of four participants completed one image-sorting task using the same images and categories. Participants were instructed to work together to reach a consensus on the grouping of each image.
@@ 3.3.1 Primary Task
Participants sorted 28 images from the OASIS dataset [31], which includes 900 validated images rated by 822 participants for pleasantness and arousal. Images were selected to span diverse emotions while avoiding distressing content. Participants sorted the selected images into one of six emotion categories randomly chosen from Russell’s circumplex model of affect [49]: angry, bored, relaxed, tense, pleased, or frustrated. Image sorting tasks have been shown to foster decision-making, communication, and social coordination by building shared mental models and group alignment [10]. We focus on this collaborative task marked by asynchronous, flexible participation, where participants can contribute and modify inputs independently. This repeated image-sorting task, involving open-ended discussions on the emotions elicited by each image, allows us to observe a dynamic, iterative collaborative process among the group, where participants achieve a shared goal through incremental steps and consensus. Each group of four participants completed one image-sorting task on the same 28 images and categories. Participants were instructed to work together as a group to reach a consensus on the label of each image, with no time limit for completing the task, allowing participants to engage in deliberate discussion and negotiation. The labels are not placed in a fixed position, allowing participants to organize and use the room space however they want. The task ends once they inform the researcher that they agree with the image groupings.
@@ 3.3.2 Virtual Scene and Cues
As shown in What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality (left, headset’s first-person perspective), all 28 virtual images are scattered around the room, and all six emotion category labels (gray virtual plates) are pasted a little higher than where the images are scattered. Participants can view these images and labels at all times via their headsets. For participants, the cue to start the interaction is not defined by the researcher but decided by each participant which image they want to discuss with the group to sort. This lack of structure in cues is by design, as our goal is to observe open and free collaborations without participants taking turns or being directed by the flow of the virtual scene designed by the researcher. The participants are assumed to take the cue for virtual interaction from the other three group members as shown in What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality (right), where all four participants are physically close, probably examining the same image and discussing the final label.
@@ 3.3.3 Interaction and Feedback
To sort an image, the participants were asked to physically move the virtual image near the virtual label. Once the image is pasted close to the label, the image is recorded as sorted. Participants used a point-and-drag near-interaction motion with the grip buttons on their left or right controllers to move an image to the corresponding label. They pointed at an image, pressed and held the grip button to drag it, and released it to place it. The image followed the controller’s pointer and locked into position when released. Participants could only grab objects within reach and were instructed to release the grip button to secure an image in place. Only one participant could move an image at a time, but multiple images could be held simultaneously. This interaction mimicked real-world object placement and provided immediate feedback once placed.
@@ 3.4 Measurements
This section outlines the measures we used to capture group behavior in the image sorting task, categorized into sensor level, task-related performance metrics, and subjective measures. At the sensor level, we collected data from the headset using custom scripts. We recorded the audio signal from the microphone, positions for location, and eye-tracking data to capture the data for conversation, proximity, and shared attention as described in §2.1.2.
@@
Finally, we collect post-task subjective measures via surveys after the image sorting task, such as presence with the IPQ [51] and PQ [68] questionnaires, cognitive load through NASA-Task Load Index (NASA-TLX) [23]222We are aware of the criticism surrounding the use of NASA-TLX. We applied it to measure “perceived” cognitive workload, rather than actual mental load [35]., and perception of the group behavior through a custom-designed questionnaire. The PQ evaluates factors such as the possibility to act and examine, realism, self-evaluation, and interface quality. The IPQ measures spatial and general presence, realism, and involvement. The presence scores are derived from 33 items (14 IPQ and 19 PQ, the cognitive load score is derived from 5 items from NASA-TLX, and group behavior from 5 items from our custom-designed survey on a 7-point scale. We developed a custom questionnaire to measure participants’ perspectives of how their group interacted, as shown in Table 2. The proposed group behavior characterization questionnaire assesses key aspects such as contribution awareness, shared attention, proximity impact, conversational support, and overall group collaboration. Each question is informed by established research to ensure relevance to our study’s context.
@@ 3.5 Study Procedure
Upon arrival, participants received a detailed information sheet outlining procedures, data collection, and privacy. The researcher also provided a verbal briefing, covering headset interactions, controller gestures, and visual stimuli details such as their color, shape, duration, cues, and feedback mechanism for the image-sorting task. Participants were given ample time to consider their participation in the study and were asked for their verbal consent. Participants then completed a demographic questionnaire covering age, gender, tech familiarity, and familiarity with group members.
MR headsets were then distributed to the participants, and they were instructed to calibrate the focus and fit of the headset for maximum comfort. Before the main task, participants completed a tutorial application with two images and two categories not included in the main task to prevent learning effects. This tutorial task aimed to familiarize them with the task interactions in terms of gestures and get them comfortable with using the point-and-drag interaction from the controller. Following this, the group proceeded with the main image-sorting task. They were informed that there was no time limit for the task and that the main requirement was for them to reach a consensus on the image sorting for the task to end. To encourage a more natural and unconstrained group behavior, participants were not informed that they were being timed or evaluated on their accuracy.
Upon task completion, participants filled out a post-task questionnaire on cognitive load, presence, and perceived collaboration. The time it took for each group to complete the task varied; however, the total duration of the session, including consent, briefing, training, calibration, task, and surveys, lasted under an hour.
Metrics: Mean (), Standard Deviation (), Standard Error (SE), 95% Confidence Interval (CI), 5th/95th Percentiles (P5/P95), Min, and Max.
@@
4 Results
4.1 Perceived Group Behavior & Tasks Summary
We summarize the descriptive statistics illustrating participants’ subjective experiences with response variability in Table 3. Metrics such as mean () and standard deviation () PQ () and IPQ () indicate moderate presence levels. We also reported sub-scales: realism (PQ-REAL, IPQ-REAL), possibility to act (ACT), interface quality (IFQUAL), possibility to examine (EXAM), self-evaluation of performance (EVAL), involvement (INV), general presence (GP), and spatial presence (SP). The PQ-REAL subscale (), suggests moderate realism, while the ACT subscale shows a high , indicating strong perceived action capability. INV subscale variability () highlights differing engagement levels. SP () reflects strong spatial awareness. The NASA TLX score () points to a low perceived workload. Group cohesion () and collaboration scored high (), while group proximity varied (), indicating differing closeness perceptions.
Next, we present two complementary sets of group-level metrics. A summary of behavioral interaction statistics captured through system instrumentation in Table 4. These include the number of images grabbed per participant, total grabs, label overrides, label changes, and task completion time. These metrics reflect the group’s engagement with the task, coordination complexity, and task performance. For example, Group 8 had a notably high number of total grabs (436) and the longest task duration (994s), suggesting prolonged deliberation or difficulty reaching consensus, whereas Group 5 completed the task most quickly (209s), with lower override activity, possibly reflecting more streamlined decision-making or higher initial agreement.
@@ 4.2 Collaboration via Sensor-based Indicators
This section discusses the participants’ reflections on collaboration, alongside behavioral indicators obtained from sensors and system logs, at both individual and group scales. Findings for shared gaze overlap frequency (indicating visual focus on the same virtual image), proximity duration (measuring time spent in physical proximity during interaction), and speaking proportion and variance (representing turn-taking and conversational balance via total speaking time, frequency, and variance in participation) can be seen in Table 6. Normality checks for subjective collaboration scores were conducted using the Shapiro-Wilk test ( at the group level; - at the individual level), determining the choice of Spearman or Pearson test based on sensor metric normality. Highlighted table rows indicate statistically significant correlations (), offering quantitative insight into how subjective assessments correlate with interaction behavior patterns.
@@ 4.4 Experience via Sensor-based Indicators
We first present the correlations between presence and cognitive load scores and sensor-derived indicators of shared attention, proximity, and interaction balance in Table 8. Sensor-based metrics were drawn from the same behavioral indicators introduced earlier, capturing visual coordination, physical closeness, and contribution variability. Presence was not significantly associated with shared attention metrics (shared attention percentage: , ), proximity (mean pairwise distance: , ), or interaction balance (variance in image grabs: , ). This suggests that the individual sense of presence was not directly mirrored in the observable group interaction patterns. For cognitive load, however, we found a positive association with shared gaze overlap at the individual level (, ), indicating that users reporting higher mental effort tended to participate in more episodes of shared attention. Other indicators, including proximity and interaction balance, did not show significant relationships with cognitive load (, ). These results suggest that internal experience is partially reflected in the observable sensor-based group metrics. In particular, shared visual focus may demand or reinforce cognitive effort. At the same time, spatial positioning and contribution balance may reflect other aspects of collaboration beyond individual mental workload or presence.
@@ 5 Discussion
This section analyzes findings in the context of our S2S framework by assessing objective behavioral indicators, subjective experiences, and group reflections. We evaluate how these signals match participants’ reported collaboration experiences and task results, revealing insights into how collaboration is both enacted and experienced. Significant associations were found between sensed behavior and reported experiences, but only at the group level did they become truly robust. We distinguish between individual and group insights to evaluate the alignment’s strength and origin. Shared gaze frequency was a consistent indicator of perceived collaboration and shared attention at both analysis levels. Group-level correlations were particularly robust, highlighting shared visual focus as both an individual and collective phenomenon, aligning with Bai et al. [4] who showed that combining gaze and gesture cues enhances co-presence and collaboration effectiveness. In contrast, conversational support exhibited weaker correlations. A significant relationship between speaking balance and perceived support was only identified at the group level, implying that individual perceptions may not align with supportive speaking dynamics, but group-level participation does indicate communicative balance (supporting established findings in [60]). We accept H1 and H1.1, as multiple gaze-based features correlated with perceived collaboration and shared attention. We partially confirm H1.2, as conversational dynamics were significant solely at the group level.
@@
Behavioral indicators, particularly shared gaze frequency and proximity duration, demonstrated strong predictive ability for task completion time, especially at the group level. Groups that engaged in mutual observation and remained close completed tasks more rapidly, underscoring the importance of co-orientation in effective collaboration, as shown by our study’s metrics. Interaction balance, such as variance in image grabs or label overrides, was also linked with efficiency. Notably, these patterns persisted at both individual and group scales, reinforcing the role of participation equity in enhancing fluency [17]. We accept H2, H2.1, and H2.2. Group-level behavior metrics, notably shared attention and proximity, predicted performance, and interaction balance indicated collaborative fluency across both levels.
@@
Sensory behavioral features were weakly associated with subjective states such as the presence or cognitive load. A near-significant individual-level link was observed between cognitive load and shared gaze frequency, but it lacked strength and consistency at the group level. Presence did not correspond clearly with proximity or gaze patterns, suggesting a lack of visible manifestation of internal states in this context. Likewise, no significant relationship emerged between cognitive load and participation balance. H3 is partially accepted due to some individual-level signal, but not consistently. We reject H3.1, H3.2, and H3.3 due to a lack of strong evidence connecting presence or cognitive load to group-level behaviors. Although disappointing at first glance, this result replicates concerns by others [9, 25] that objective signals often fail to capture subjective experience unless paired with richer data sources.